The domain of the function define by f(x) = √[Inx/ (Inx-1)] is?


1
Expert's answer
2021-08-24T16:22:53-0400
"\begin{cases}
 \dfrac{\ln x}{\ln x-1}\geq 0 \\
 x>0 \\
\ln x\not=1
\end{cases}"

"\begin{cases}
 \ln x\leq0\ or\ \ln x> 1 \\
 x>0 \\

\end{cases}"

Domain: "(0, 1]\cup(e, \infin)"